OpenStudy (anonymous):

Stuck... Whats the next step? Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Step 4: ?

OpenStudy (azureilai):

I think it would be x=-4+2 and x = 4+2 because it is an absolute value equation, it works alot like a quadratic when you square root it. since -2-2 = -4which becomes 4 in an absolute value equation, and 6-2 =4

OpenStudy (anonymous):

x + 2 = -4 -x - 2 = 4 x + 2 = 4 x - 2 = -4 Where the choices @Azureilai

OpenStudy (azureilai):

the last one will work, since x=-2, and if you plug it in, the equation is true. |(-2) - 2| = 7 - 3 |-4| = 4 4=4